Construct a simulated 1H NMR spectrum, including p

Construct a simulated 1H NMR spectrum, including proton integrations, for CH3OC(CH2OCH3)3 (see Hint). Drag the appropriate splitting patterns to the approximate chemical shift positions; place the integration values in the small bins above the associated chemical shift. Splitting patterns and integrations may be used more than once, or not at all, as needed. Likewise, some bins might remain blank. Note that peak heights are arbitrary and do not indicate proton integrations.
Answer #1
Concepts and reason

The proton NMR spectrum:

• The proton NMR spectroscopy is one of the techniques that are useful to predict the structure of the compound.

• In 1HNMR^{\rm{1}}{\rm{H NMR}} spectroscopy, peaks are observed at the point where the wavelengths of proton nuclei are matched to substance nuclei wavelengths.

• In the same manner, there are other spectroscopies like 13CNMR^{{\rm{13}}}{\rm{C NMR}} , IR and mass spectroscopy that are present.

Fundamentals

1HNMR^{\rm{1}}{\rm{H NMR}} Spectroscopy

In proton NMR spectroscopy, the number of peaks’ value depends on the number of hydrogen atoms attached to that carbon position.

numberofline=2nI+1I=1/2forprotonn=numberofhydrogenatomsatadjecentcarbonatom\begin{array}{l}\\{\rm{number of line = 2nI + 1}}\\\\{\rm{I = 1/2 for proton}}\\\\{\rm{n = number of hydrogen atoms at adjecent carbon atom}}\\\end{array}

Given structure is drawn below.

HаС—
-CH3З
Н2
-с
-CH3
HаС-
-o-
Нас—
о
-CHз

It is clear that three different types of protons are present in the above structure. Nine hydrogen atoms are having identical behavior (methyl groups), six hydrogen atoms are having identical behavior, and three hydrogen atoms have identical behavior (methyl group adjacent to oxygen).

Three singlet peaks are observed in proton NMR spectra because none of the types have adjacent protons.

Singlet
Singlet
Singlet
In between 3 and 4
9H
Around 4ppm
Around 3ppm
бн
зн

The proton NMR spectrum is drawn below.

9H
6H
3H
ppm

Ans:

9H
6H
3H
ppm